# Topological constraints on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v2

@@ 2 Symplectic field theory for exact Lagrangian cobordisms
We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $2c_{1}(T\bar{P})=0$. Let us consider
$\begin{array}[]{ll}Y\coloneqq P\times\mathbb{R},&\alpha\coloneqq dz-\lambda,\end{array}$
where $z$ is the coordinate of $\mathbb{R}$. Then, $\alpha$ is a contact form on $Y$ and the Reeb vector field on $Y$ with respect to $\alpha$ agrees with $\partial_{z}$. Let $\pi_{P}\colon Y\to P$ denote the projection to $P$.
The arguments from Section 2.1 to Section 2.4 are parallel to those in [16, Section 3], except the difference of the coefficient rings. In Section 2.5, we will discuss a particular case related to clean Lagrangian intersections.
@@ 2.1 Preliminaries
$\psi_{0}$ $\colon[0,\infty)\times[0,1]\to D_{m+1},$
$\psi_{k}$ $\colon(-\infty,0]\times[0,1]\to D_{m+1}\text{ for }k=1,\dots,m,$
with respect to $j_{0}$ and $j_{\kappa}$ such that $\lim_{s\to\infty}\psi_{0}(s,t)=p_{0}$ and $\lim_{s\to-\infty}\psi_{k}(s,t)=p_{k}$ uniformly on $t\in[0,1]$.
@@
Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. We denote the set of Reeb chords of $\Lambda$ by
@@
Hereafter, for any Reeb chord $(a\in[0,T]\to Y)\in\mathcal{R}(\Lambda)$, we write $T=\int a^{*}\alpha$ as $T_{a}$. For each $a\in\mathcal{R}(\Lambda)$, we fix a continuous map $w_{a}\colon\overline{D_{1}}\to Y$ such that $w_{a}(\partial D_{1})\subset\Lambda$ and
$\lim_{s\to\infty}w_{a}\circ\psi_{0}(s,t)=a(T_{a}t)$
for every $t\in[0,1]$.
@@ 2.2 Moduli space of pseudo-holomorphic curves with boundary in an exact Lagrangian cobordism
  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and its restriction
    $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}\colon(-\infty,r_{-}]\times\Lambda_{-}\to\mathbb{R}$
    is a constant function. Here, $r$ is the $\mathbb{R}$-coordinate of $\mathbb{R}\times Y$.
@@
$\begin{split}i_{+}&\colon(Y,\Lambda_{+})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{+},x),\\ i_{-}&\colon(Y,\Lambda_{-})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{-},x),\end{split}$ (1)
@@
For any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$, we define
$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$
to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to\mathbb{R}\times Y$ such that:
@@
    $\lim_{s\to\infty}\tau_{-s}\circ u\circ\psi_{0}$ $=(s_{0},a(T_{a}(1-t))),$
    $\lim_{s\to-\infty}\tau_{s}\circ u\circ\psi_{k}$ $=(s_{k},b_{k}(T_{b_{k}}(1-t)))\text{ for }k=1,\dots,m,$
    $C^{\infty}$-uniformly on $[0,1]$.
@@ 2.3 Chekanov-Eliashberg DGA over $\mathbb{Z}[H_{2}(Y,\Lambda)]$
The second condition is satisfied for generic $\Lambda$. Under the second condition, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as in [16, Section 3.1]. We then define
$|a|\coloneqq\mu(a)-1.$
We also assume that $\Lambda$ is admissible in the sense of the definition in [5, Section 2.3]. Let us fix a spin structure $\mathfrak{s}$ on $\Lambda$.
We take an almost complex structure $J$ on $P$ as in [5, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$ such that $(b_{1},\dots,b_{m})\neq(a)$: The moduli space $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$. Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold. (For the relation between $J$-holomorphic disks in $P$ defined as in [5, Section 2.3] and $\widetilde{J}-$holomorphic maps in $\mathbb{R}\times Y$ defined as in Section 2.2, see [2, Theorem 2.1] and [16, Appendix A.4].)
In addition, $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$. See [16, Definition A.4] for details which follow the conventions in [4] related to orientations.
Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose grading is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$. A derivation over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of degree 1-1
$\partial_{J,\mathfrak{s}}\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*-1}(\Lambda)$
is defined for every $a\in\mathcal{R}(\Lambda)$ by
@@ 2.4 DGA map associated to an exact Lagrangian cobordism
We consider an exact Lagrangian cobordism $L$ from $\Lambda_{-}$ to $\Lambda_{+}$ as in Section 2.2 such that the Maslov class of $L$ vanishes and both $\Lambda_{+}$ and $\Lambda_{-}$ are connected. We assume that $L$ is oriented and has a spin structure $\mathfrak{s}_{L}$. Then, on positive end $\Lambda_{+}$ (resp. the negative end $\Lambda_{-}$) of $L$, an orientation and a spin structure $\mathfrak{s}_{+}$ (resp. $\mathfrak{s}_{-}$) are induced. See [16, Remark 3.6] for their precise descriptions.
Associated to $(\Lambda_{+},\mathfrak{s}_{+})$ and $(\Lambda_{-},\mathfrak{s}_{-})$, the Chekanov-Eliashberg DGAs
$\begin{array}[]{cc}(\mathcal{A}_{*}(\Lambda_{+}),\partial_{J,\mathfrak{s}_{+}}),&(\mathcal{A}_{*}(\Lambda_{-}),\partial_{J,\mathfrak{s}_{-}})\end{array}$
are defined as in Section 2.3. The maps $i_{+}$ and $i_{-}$ of (1) induce ring homomorphisms
@@
By [7, Lemma 3.8], the moduli space $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is cut out transversely for generic $J$ and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)$ for any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$. By the Gromov compactness, if $|b_{1}|+\dots+|b_{m}|=|a|$, $\coprod_{C\in H_{2}(\mathbb{R}\times Y,L)}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold. In addition, it has an orientation depending on $\mathfrak{s}_{L}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda_{+}$ and $\Lambda_{-}$. See [16, Section A.3].
We refer to the proof of [16, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings. We modify the arguments in [16, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings. (As referred in [16, Appendix A], see also [10, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [10] related to orientations are different from those in [4, 16].) Then, we obtain the following theorem.
@@ 2.5.1 Chekanov-Eliashberg DGA of the unit conormal bundle of a knot in $\mathbb{R}^{3}$
Let $(q_{1},q_{2},q_{3})$ be the coordinate of $\mathbb{R}^{3}$. The manifold $\mathbb{R}^{3}$ is equipped with a standard Riemannian metric $\langle\cdot,\cdot\rangle\coloneqq\sum_{i=1}^{3}dq_{i}\otimes dq_{i}$ and an orientation such that $dq_{1}\wedge dq_{2}\wedge dq_{3}$ is a positive volume form. A coordinate $((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))$ of $T^{*}\mathbb{R}^{3}$ is defined via the diffeomorphism
$\mathbb{R}^{3}\times\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto\sum_{i=1}^{3}p_{i}(dq_{i})_{(q_{1},q_{2},q_{3})}.$
For any $r>0$, we denote $D^{*}_{r}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|\leq r\}$. The canonical Liouville form on $T^{*}\mathbb{R}^{3}$ is given by $\lambda_{\mathbb{R}^{3}}=\sum_{i=1}^{3}p_{i}dq_{i}$.
We consider the unit cotangent bundle of $\mathbb{R}^{3}$
$S^{*}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|=1\}.$
It is endowed with a contact form $\alpha_{\mathbb{R}^{3}}\coloneqq\left.\lambda_{\mathbb{R}^{3}}\right|_{S^{*}\mathbb{R}^{3}}$. Note that there exists a diffeomorphism
@@
We focus on a special kind of Legendrian submanifolds. For any knot $K$ in $\mathbb{R}^{3}$, recall its conormal bundle $L_{K}$ defined in Section 1. Then, the unit conormal bundle of $K$ is defined by
@@
Let us prepare several notations. We identify a disk bundle of $L_{K}$ with a tubular neighborhood of $K$ in $\mathbb{R}^{3}$ via the map $L_{K}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\mathbb{R}^{3}\colon(q,p)\mapsto q+p$ for sufficiently small $\varepsilon>0$. Then, an orientation of $L_{K}$ is determined from the orientation of $\mathbb{R}^{3}$. We also fix a spin structure $\mathfrak{s}_{L_{K}}$ given by a trivial $\mathrm{Spin}(3)$-bundle on $L_{K}$. As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [16, Remark 3.6].
@@
We choose a non-vanishing vector field $V$ on $K$. Then, $\Lambda_{K}$ is bounded by a compact orientable manifold
$\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ p(V_{q})\geq 0\}.$
Therefore, the homology class $\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
@@
We denote $\lambda\coloneqq e^{\widehat{l}}$, $\mu\coloneqq e^{\widehat{m}}$ and $U\coloneqq e^{\widehat{s}}$, then
$\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})]=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}].$
From the arguments in Section 2.3, the Chekanov-Eliashberg DGA $(\mathcal{A}_{*}(\Lambda_{K}),\partial_{J,\mathfrak{s}_{K}})$ for $(\Lambda_{K},\mathfrak{s}_{K})$ is defined with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.

### v3

@@ 2 Symplectic field theory for exact Lagrangian cobordisms
We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $2c_{1}(T\bar{P})=0$. Let us consider $Y\coloneqq P\times\mathbb{R}$ and $\alpha\coloneqq dz-\lambda$, where $z$ is the coordinate of $\mathbb{R}$. Then, $\alpha$ is a contact form on $Y$ and the Reeb vector field on $Y$ with respect to $\alpha$ agrees with $\partial_{z}$. Let $\pi_{P}\colon Y\to P$ denote the projection to $P$.
The arguments from Section 2.1 to Section 2.4 are parallel to those in [15, Section 3], except the difference of the coefficient rings. In Section 2.5, we will discuss a particular case related to clean Lagrangian intersections.
@@ 2.1 Preliminaries
$\begin{array}[]{cc}\psi_{0}\colon[0,\infty)\times[0,1]\to D_{m+1},&\psi_{k}\colon(-\infty,0]\times[0,1]\to D_{m+1}\end{array}$
for $k\in\{1,\dots,m\}$ with respect to $j_{0}$ and $j_{\kappa}$ such that $\lim_{s\to\infty}\psi_{0}(s,t)=p_{0}$ and $\lim_{s\to-\infty}\psi_{k}(s,t)=p_{k}$ uniformly on $t\in[0,1]$.
@@
Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. Denote the set of its Reeb chords by
@@
Hereafter, for any Reeb chord $(a\in[0,T]\to Y)\in\mathcal{R}(\Lambda)$, we write $T=\int a^{*}\alpha$ as $T_{a}$. For each $a\in\mathcal{R}(\Lambda)$, we fix a continuous map $w_{a}\colon\overline{D_{1}}\to Y$ such that $w_{a}(\partial D_{1})\subset\Lambda$ and $\lim_{s\to\infty}w_{a}\circ\psi_{0}(s,t)=a(T_{a}t)$ for every $t\in[0,1]$.
@@ 2.2 Moduli space of pseudo-holomorphic curves with boundary in an exact Lagrangian cobordism
  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and $f$ is constant when restricted on $(-\infty,r_{-}]\times\Lambda_{-}$. Here, $r$ is the $\mathbb{R}$-coordinate of $\mathbb{R}\times Y$.
@@
$\begin{array}[]{cc}i_{+}\colon(Y,\Lambda_{+})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{+},x),&i_{-}\colon(Y,\Lambda_{-})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{-},x),\end{array}$ (2)
@@
For any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$, we define $\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$ to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to\mathbb{R}\times Y$ such that:
@@
    $\begin{array}[]{cc}\lim_{s\to\infty}\tau_{-s}\circ u\circ\psi_{0}=(s_{0},a(T_{a}(1-t))),&\lim_{s\to-\infty}\tau_{s}\circ u\circ\psi_{k}=(s_{k},b_{k}(T_{b_{k}}(1-t)))\end{array}$
    for $k\in\{1,\dots,m\}$, $C^{\infty}$-uniformly on $[0,1]$.
@@ 2.3 Chekanov-Eliashberg DGA over $\mathbb{Z}[H_{2}(Y,\Lambda)]$
The second condition is satisfied for generic $\Lambda$. Under the second condition, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as in [15, Section 3.1]. We then define $|a|\coloneqq\mu(a)-1$. We also assume that $\Lambda$ is admissible in the sense of the definition in [5, Section 2.3]. Let us fix a spin structure $\mathfrak{s}$ on $\Lambda$.
We take an almost complex structure $J$ on $P$ as in [5, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$ such that $(b_{1},\dots,b_{m})\neq(a)$: The moduli space $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$. Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold. (For the relation between $J$-holomorphic disks in $P$ defined as in [5, Section 2.3] and $\widetilde{J}-$holomorphic maps in $\mathbb{R}\times Y$ defined as in Section 2.2, see [2, Theorem 2.1] and [15, Appendix A.4].)
In addition, $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$. See [15, Definition A.4] for details which follow the conventions in [4] related to orientations.
Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose grading is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$. A derivation over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ $\partial_{J,\mathfrak{s}}\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*-1}(\Lambda)$ is defined for every $a\in\mathcal{R}(\Lambda)$ by
@@ 2.4 DGA map associated to an exact Lagrangian cobordism
We consider an exact Lagrangian cobordism $L$ from $\Lambda_{-}$ to $\Lambda_{+}$ as in Section 2.2 such that the Maslov class of $L$ vanishes and both $\Lambda_{+}$ and $\Lambda_{-}$ are connected. We assume that $L$ is oriented and has a spin structure $\mathfrak{s}_{L}$. Then, on positive end $\Lambda_{+}$ (resp. the negative end $\Lambda_{-}$) of $L$, an orientation and a spin structure $\mathfrak{s}_{+}$ (resp. $\mathfrak{s}_{-}$) are induced. See [15, Remark 3.6] for their precise descriptions.
Associated to $(\Lambda_{+},\mathfrak{s}_{+})$ and $(\Lambda_{-},\mathfrak{s}_{-})$, the Chekanov-Eliashberg DGAs $(\mathcal{A}_{*}(\Lambda_{+}),\partial_{J,\mathfrak{s}_{+}})$ and $(\mathcal{A}_{*}(\Lambda_{-}),\partial_{J,\mathfrak{s}_{-}})$ are defined as in Section 2.3. The maps $i_{+}$ and $i_{-}$ of (2) induce ring homomorphisms
@@
By [7, Lemma 3.8], the moduli space $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is cut out transversely for generic $J$ and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)$ for any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$. By the Gromov compactness, if $|b_{1}|+\dots+|b_{m}|=|a|$, $\coprod_{C\in H_{2}(\mathbb{R}\times Y,L)}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold. In addition, it has an orientation depending on $\mathfrak{s}_{L}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda_{+}$ and $\Lambda_{-}$. See [15, Section A.3].
We refer to the proof of [15, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings. We modify the arguments in [15, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings. (As referred in [15, Appendix A], see also [9, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [9] related to orientations are different from those in [4, 15].) Then, we obtain the following theorem.
@@ 2.5.1 Chekanov-Eliashberg DGA of the unit conormal bundle of a knot in $\mathbb{R}^{3}$
Let $(q_{1},q_{2},q_{3})$ be the coordinate of $\mathbb{R}^{3}$. The manifold $\mathbb{R}^{3}$ is equipped with a standard Riemannian metric $\langle\cdot,\cdot\rangle\coloneqq\sum_{i=1}^{3}dq_{i}\otimes dq_{i}$ and an orientation such that $dq_{1}\wedge dq_{2}\wedge dq_{3}$ is a positive volume form. Let $(p_{1},\dots,p_{3})$ be the coordinate of the fiber. For any $r>0$, we denote $D^{*}_{r}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|\leq r\}$. The canonical Liouville form on $T^{*}\mathbb{R}^{3}$ is given by $\lambda_{\mathbb{R}^{3}}=\sum_{i=1}^{3}p_{i}dq_{i}$.
We consider the unit cotangent bundle $S^{*}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|=1\}$. It is endowed with a contact form $\alpha_{\mathbb{R}^{3}}\coloneqq\left.\lambda_{\mathbb{R}^{3}}\right|_{S^{*}\mathbb{R}^{3}}$. Note that there exists a diffeomorphism
@@
We focus on a special kind of Legendrian submanifolds. For any knot $K$ in $\mathbb{R}^{3}$, let $L_{K}$ be its conormal bundle. Then, the unit conormal bundle of $K$ is defined by
@@
Let us prepare several notations. We identify a disk bundle of $L_{K}$ with a tubular neighborhood of $K$ in $\mathbb{R}^{3}$ via the map $L_{K}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\mathbb{R}^{3}\colon(q,p)\mapsto q+p$ for sufficiently small $\varepsilon>0$. Then, an orientation of $L_{K}$ is determined from the orientation of $\mathbb{R}^{3}$. We also fix a spin structure $\mathfrak{s}_{L_{K}}$ given by a trivial $\mathrm{Spin}(3)$-bundle on $L_{K}$. As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [15, Remark 3.6].
@@
We choose a non-vanishing vector field $V$ on $K$. Then, $\Lambda_{K}$ is bounded by a compact orientable manifold $\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ p(V_{q})\geq 0\}$. Therefore, the homology class $\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
@@
We denote $\lambda\coloneqq e^{\widehat{l}}$, $\mu\coloneqq e^{\widehat{m}}$ and $U\coloneqq e^{\widehat{s}}$, then $\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})]=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. From the arguments in Section 2.3, the Chekanov-Eliashberg DGA $(\mathcal{A}_{*}(\Lambda_{K}),\partial_{J,\mathfrak{s}_{K}})$ for $(\Lambda_{K},\mathfrak{s}_{K})$ is defined with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.